Use the Lagrange multipliers to solve the following:

Maximize f(x,y,z) = 4x + 2y + z

subject to x2 + y + z2 = 1


1
Expert's answer
2022-05-10T23:39:18-0400

Let’s define "g(x, y, z) = x^ 2 +y +z^ 2 ," so the problem is to find the maximum of "f(x, y, z)" subject to the constraint "g(x, y, z) = 1." We have 


"\nabla f=\lambda \nabla g"

"\langle4,2,1\rangle=\lambda\langle2x,1,2z\rangle"

Reading this component by component and including the restriction we get the system of equations 


"4=2\lambda x"

"2=\lambda"

"1=2\lambda z"

"x^ 2 +y+z^ 2=1"

Then "x=1, z=\dfrac{1}{4}, \lambda=2."


"1+y+\dfrac{1}{16}=1"


"y=-\dfrac{1}{16}"

The maximum of "f" occurs when "x=1, y=-\dfrac{1}{16}, z=\dfrac{1}{4}."

"f(1,-\dfrac{1}{16},\dfrac{1}{4}) = 4(1) + 2(-\dfrac{1}{16}) + \dfrac{1}{4}=\dfrac{33}{8}"